### v1

# Sharing is CAIRing: Characterizing Principles and Assessing Properties of Universal Privacy Evaluation for Synthetic Tabular Data

Abstract
Data sharing is a necessity for innovative progress in many domains, especially in healthcare. However, the ability to share data is hindered by regulations protecting the privacy of natural persons. Synthetic tabular data provide a promising solution to address data sharing difficulties but does not inherently guarantee privacy. Still, there is a lack of agreement on appropriate methods for assessing the privacy-preserving capabilities of synthetic data, making it difficult to compare results across studies. To the best of our knowledge, this is the first work to identify properties that constitute good universal privacy evaluation metrics for synthetic tabular data. The goal of such metrics is to enable comparability across studies and to allow non-technical stakeholders to understand how privacy is protected. We identify four principles for the assessment of metrics: Comparability, Applicability, Interpretability, and Representativeness (CAIR). To quantify and rank the degree to which evaluation metrics conform to the CAIR principles, we design a rubric using a scale of 1-4. Each of the four properties is scored on four parameters, yielding 16 total dimensions. We study the applicability and usefulness of the CAIR principles and rubric by assessing a selection of metrics popular in other studies. The results provide granular insights into the strengths and weaknesses of existing metrics that not only rank the metrics but highlight areas of potential improvements. We expect that the CAIR principles will foster agreement among researchers and organizations on which universal privacy evaluation metrics are appropriate for synthetic tabular data.
1 Introduction
The volume of data collected and produced continues to increase as our society moves toward a higher degree of digitalization, and personal and health data are no exception. Regulatory restrictions on data sharing, such as HIPAA [15] and GDPR [7], have been implemented to protect individuals from privacy violations. The downside of imposing restrictions on data sharing is the substantial limitations on innovation and research based on those data. Consequently, methods that allow data sharing under the regulations are a necessity to catalyze data-sharing capabilities and privacy improvements. Privacy-preserving synthetic data generation is a promising attempt to solve this problem by generating simulated records with the same statistical properties as the true data [18, 17]. However, as synthetic data generation is still maturing for tabular data, an abundance of evaluation metrics are applied in various studies. Consequently, comparing the evaluation results of the generated data is a highly challenging task, which slows innovation in generation methods [2, 8, 11]. Therefore, standardized approaches are needed to evaluate the quality of synthetic data. In addition, more regulations are in the works as of writing this paper, namely the European Union Artificial Intelligence Act [6] which moves towards a more secure, fair, and private use of artificial intelligence (AI). Therefore, privacy evaluation is important when dealing with synthetic data. Many studies claim that fully synthetic data are inherently private [11, 14] due to the disconnect between true and synthetic data. However, it has been shown that synthetic data can be successfully attacked and therefore require a thorough evaluation of privacy [12, 4].
When we combine the knowledge that 1) synthetic data are promising, 2) there is no consensus for evaluation metrics, 3) regulations are focusing more on privacy in AI, and 4) synthetic data are not inherently private; it is clear that we need to define how we best assess the privacy levels provided by synthetic data. The goal is to allow for comparability across works in a way that allows non-technical stakeholders, such as regulators and natural persons, to understand how privacy is protected.
Therefore, the aim of this paper is to define principles for good universal privacy evaluation metrics for synthetically generated tabular data that can aid researchers and regulators in creating and selecting appropriate evaluation methods. We have created a scoring system that uses a rubric to rank existing evaluation metrics in order to achieve that goal. In this work, we use the concept of privacy metric and privacy measure synonymously for any function, method, algorithm, or framework that is used to evaluate the privacy level provided by synthetically generating data. To the best of our knowledge, this is the first attempt to define good practices with a scoring system for the evaluation of the privacy of synthetic tabular data.
In the remainder, we present related work (Section 2), followed by a presentation of the criteria and rubric developed in this work (Section 3). Then we detail all dimensions of the rubric (Section 4) and apply them to a selection of existing privacy evaluation metrics (Section 5). Lastly we discuss the strengths, limitations, and significance of our work (Section 6).
2 Related Work
One of the primary motivations for creating synthetic data is to introduce privacy, but the evaluation of utility and realism has been the main focus of many researchers [14]. The same focus on utility is also present when defining universally applicable evaluation metrics, with most research revolving around finding suitable utility or realism metrics. El Emam et al. [5] attempt to quantify how well six different utility metrics can rank the performance of synthetic data generation methods according to a logistic regression prediction. They empirically evaluate 30 different datasets using three SDG methods. The baseline against which they test is the difference between prediction performance for the true and synthetic data using the area under the receiver operating characteristic curve and the area under the precision-recall curve. Using the Page test, they test whether the utility metrics agree with the baseline. This approach enables empirical quantification of the metrics’ performances but is limited to the single scenario of using logistic regression. Dankar et al. [3] identify four classes of utility metrics for which they select a representative utility metric that they test. Ultimately, they assess whether one of the metrics can be used for overall utility assessment by investigating the correlations between the selected metrics. The authors identify the Hellinger distance as somewhat correlated with two of the three metrics and conclude that it can be used as a standardized utility metric. The study focuses on a limited set of criteria for streamlining evaluation of utility and does not consider elements such as attribute heterogeneity. Hernandez et al. [10, 11] identify the issue of comparing evaluation results across studies. Therefore, they opt to use a proxy scale consisting of “excellent”, “good”, and “poor” and define the criteria to fall into each level for all included evaluation metrics. The downside of using this scale is the limited granularity for interpretation. Furthermore, not all metrics fit naturally into the three levels. For example, they distribute the result for membership inference attacks (MIAs) [18] into “excellent” if privacy is preserved and “good” otherwise. Consequently, even though that method is a step in the right direction for comparability, it is non-exhaustive and provides a limited proxy for evaluation.
3 Method
Assessing the quality of complex work such as privacy evaluation metrics can be a challenging task to perform reproducibly and objectively. Therefore, we break down the assessment into smaller dimensions using a rubric. Rubrics have been successful in educational settings in terms of assessment of lecturer-student, student-lecturer, and peer assessments [1], but also provide an important assessment tool in other fields of assessment, including privacy [16]. Each dimension should be defined on a scale where the requirements to obtain a specific score are clearly defined to avoid subjectivity and bias. For this work, four general categories have been identified based on the ideal of a universal privacy evaluation metric and the stakeholders of it:
Comparability: The ideal metric is comparable across different synthetic datasets.
Applicability: The ideal metric is applicable to all datasets, data types, and SDG methods.
Interpretability: The ideal metric can be easily communicated and understood unambiguously by stakeholders.
Representativeness: The ideal metric accurately represents a realistic privacy level.
Figure 1: The four CAIR principles with their respective dimensions.
The four categories form the acronym CAIR and represent the fact that a universal metric for tabular synthetic data should be usable for all data (C and A) and trustworthy (I and R). Each category is further broken down into four dimensions, making the rubric consist of 16 total assessments (illustrated in Figure 1). Having the same number of unweighted dimensions for each category ensures equal weighting for the general categories as well. The final result is an unweighted mean of all 16 scores with the standard error. For the sake of accuracy, we allow the inclusion of .5 decimals if a metric is deemed to fall between two scores. The complete rubric is presented in Table 1 using a scale of 1-4 (poor-excellent) with a higher score preferred. Details of each dimension are presented in Section 4.
Table 1: Rubric consisting of four categories (Comparability, Applicability, Interpretability, Representativeness), each with four dimensions. Scoring is performed on a scale of 1-4 with 0.5 decimals allowed when a metric falls between two scores. The final result is the unweighted mean of the 16 scores.
Excellent 4 Good 3 Fair 2 Poor 1
Comparability
Scale
- Completely invariant to the unit scale of attributes.
- The range of the metric is not affected by the scale of the data.
- Slightly sensitive to attributes having different scales.
- The range of the metric is not affected by the scale of the data.
- Somewhat sensitive to attributes having different unit scales.
- The scale of the metric is relative to the scale of the data.
- The metric is highly sensitive to attributes having differences in unit scales.
- The scale of the metric is relative to the scale of the data.
Metric
bounds
- The metric has a clearly defined upper and lower bound.
- The metric has an upper and lower bound.
- The bounds are defined, but they change relative to the data, e.g., can be (0, 1) for one dataset and (1, 50) for another.
- The metric has either an upper or lower bound but not both.
- The metric has no upper or lower bounds.
Data type
agnostic
- Mixed data types have little to no influence on the analysis and do not significantly affect results or conclusions.
- Mixed data types introduce mild inaccuracies in the output.
- Issues can be managed through appropriate data pre-processing or analysis techniques.
- Some combinations of mixed data types have significantly adverse effects on the analysis.
- Remediations require specialized techniques or complex efforts to handle properly.
- Mixed data types severely undermine the validity and reliability of the results, making the analysis highly complicated or infeasible.
Cross-domain
relevance
- The metric shows comprehensive relevance across diverse domains.
- Accommodates privacy and domain challenges effectively.
- The metric shows moderate cross-domain relevance, but may not cover all domains equally or comprehensively.
- The metric has limited relevance across different domains.
- Useful for more than just a narrow domain.
- The metric is highly domain-specific.
- Lacks relevance outside of a narrow domain.
Applicability
Heterogeneity
- Applicable to most tabular data types.
- Can handle mixed tabular data types in the same dataset.
- Has no bias towards certain data types.
- Applicable to most tabular data types.
- Can be modified to accommodate mixed data types, but no definition is provided.
- Has a slight bias towards certain data types.
- Applicable to at least two different data types.
- Struggles with handling mixed data types with limited adaptability.
- Has some bias towards certain data type(s).
- Applicable to only one data type.
Diverse generation methods
- Applicable to all synthetic data sets regardless of generation method.
- No bias towards any methods.
- Applicable to more than one family of generative methods.
- Has some bias towards some methods.
- Is only defined for synthetic data generated using a single family of generative methods, e.g., GANs or CARTs.
- Can only be applied to synthetic data generated with one specific generative method.
Practicality
- The required runtime is reasonable. $\mathcal{O}(n^{2})$.
- The metric balances accuracy and runtime, but further improvements to efficiency can be accomplished. $\mathcal{O}(mn^{2})$ where $1<m\ll n$.
- The metric is computationally expensive. $\mathcal{O}(n^{3})$.
- Requires substantial computational resources. $\Omega(n^{3}).$
Implementation
- Easy to implement and/or an implementation is supplied.
- The metric is computed using only readily accessible information.
- Definition or documentation of the metric has a few ambiguous elements that make consistent implementation somewhat challenging.
- The data required to compute the metric exist in the dataset and do not require information from external sources such as statistics or datasets for comparisons.
- The metric poses ambiguity and some practical challenges making it difficult to apply in real-world scenarios.
- The required information might be challenging to obtain from external sources and/or to compute.
- The metric is poorly documented and contains multiple elements open for interpretation.
- The metric is very impractical to implement, and no implementation is provided.
- The required information is very difficult or impossible to obtain from external sources and/or compute.
Interpretability
Explainability
- It is easy to communicate what the metric measures.
- The metric is straightforward and minimizes unnecessary complexity.
- The metric has some complex elements that may be difficult to communicate to a layperson in terms of what is measured.
- Computation includes unintuitive or complex elements that can be challenging to communicate to a layperson.
- The metric is complex to the degree that it is unlikely that the metric can be explained to a layperson.
- The metric requires some domain knowledge to comprehend.
- The metric is highly complex and is challenging to communicate to people with a technical background.
Understandability
- The output of the metric is easy to understand for a layperson without the need for explanations.
- The output of the metric can be understood by a layperson, but might require some explanation.
- Understanding the output of the metric requires some background knowledge and is somewhat difficult to communicate to a layperson.
- The metric contains some unnecessary complexity that could be simplified.
- The output of the metric requires substantial technical background knowledge to understand and is highly challenging to communicate to a layperson.
- The metric is overly complex and difficult to grasp.
Visualization
- The metric can easily be visualized both in isolation and compared to the results of other datasets.
- Visualizations enhance the ability to accurately compare multiple results.
- The metric can easily be visualized in isolation.
- Making comparative visualizations of multiple results in a single graph can be difficult, but can enhance the ability to differentiate between results.
- The metric can be visualized, but only in isolation.
- Comparisons with the results of other datasets are highly difficult or impossible to produce in a single or multiple graphs.
- The metric is not suited for visualization.
Granularity
- The metric offers fine-grained granularity that allows for unambiguous differentiation between privacy levels.
- Consists of a single value that makes sense in isolation.
- The metric provides a moderate level of granularity, but some important distinctions might be overlooked.
- Consists of a single value that makes sense in isolation.
- The metric has limited granularity, making it difficult to distinguish between privacy levels.
- Consists of a single value, but only makes sense when compared to other datasets or generation methods.
- The metric lacks the ability to clearly differentiate between privacy levels.
- Can consist of multiple values that describe different privacy aspects.
- Might only make sense in comparison to the privacy level of other datasets, e.g., a holdout set.
Representativeness
Anomalies
- The sensitivity of all rare observations are taken into account.
- Computation of the metric includes all observations with appropriate weights.
- Rare observations are treated with greater weight than common observations.
- Computation of the metric prioritizes rare observations.
- Rare observations are included in computing the metric.
- Rare and common observations are treated equally.
- Rare observations are not considered when computing the metric.
Coverage
- The privacy result is equally accurate for all records.
- The privacy result is approximately equally accurate for all records without any systematic bias.
- The privacy result is representative for some proper subsets of the data with a systematic bias.
- The privacy result only represents a small proper subset of the data.
Reproducibility
- The metric always returns the same value for the same input data.
- The metric returns the same value for the same input data with a probability close to 1 when applied by different people.
- The metric returns almost the same value for the same input data with some probability, but the variance is somewhat consistent when applied by different people.
- An error is provided for the output.
- The metric returns significantly different results for the same input.
Precision
- The metric does not have systematic over- or underestimation of the privacy level.
- The metric slightly underestimates the privacy level but never overestimates it.
- The metric either over- or underestimates the privacy level, but not systematically.
- The metric significantly over- or underestimates the privacy level, systematically.
4 The CAIR Principles
The CAIR principles for good privacy evaluation for synthetic tabular data consist of a range of criteria that all aim to describe the ideal privacy measure. The rubric in Table 1 is limited in description depth, but should be sufficient to score any privacy metric without additional information. However, more detailed information on the dimensions is necessary to adequately argue for their inclusion.
4.1 Comparability
One of the major arguments for having standardized privacy evaluation metrics is the ability to compare privacy results between models, datasets, and projects. Comparisons enable users to select the best-performing models and measure them against the results of others, thus driving the synthetic data generation research area in a positive direction.
C1 Scale
Comparisons of privacy levels between generation processes rely on the privacy measure being on the same scale. If the scale is not identical, the resulting evaluation values do not correspond to each other, making any attempt at comparing the results futile. An argument can be made that users should normalize the data before computing the evaluation results. However, considering that synthetically generated data are primarily useful on the same scale as the true data, it is unlikely that all users will report the results on normalized data. In addition, the scale of the data can skew the results of a measure even if the metric has well-defined bounds. Consider the function $f(\mathcal{T},\mathcal{S})=\sigma(\lVert\mathcal{T}_{i}-\mathcal{S}_{j}\rVert_{2})$, where $\mathcal{T}$ and $\mathcal{S}$ are true and synthetic data respectively and $\sigma$ is the sigmoid function; then if an attribute, $a_{0}\in(0,10^{6})$ and another attribute $a_{1}\in(0,1)$, the influence $a_{0}$ has on the function will be substantially greater than that of $a_{1}$ even though the final result is in $(0,1)$.
C2 Metric Bounds
Well-defined upper and lower bounds of a metric are important for accurately comparing evaluation results. Without fixed bounds, the results from one dataset or generation method to another do not correspond, even if the value is the same. Whether the bounds are (0, 1), (-1, 1), or something else is not important, as long as the bounds are consistent.
C3 Data Type Agnostic
Heterogeneous data types are prevalent in tabular data [14]. As such, it is crucial that the heterogeneity of the data has a limited effect on the output of the metric. Ideally, the metric should be invariant to any tabular data type, but a remediation can be to perform appropriate pre-processing. If pre-processing cannot mitigate adverse effects of the mixed data types, further analysis should be conducted to identify whether a systematic bias exists towards any specific subsets of data types.
C4 Cross-Domain Relevance
Privacy-aware synthetic data generation is relevant in a wide variety of domains. Therefore, it is no surprise that many generation methods are designed specifically for their work domains to increase utility and privacy. However, methods explicitly applied to a niche domain might be useful for research in unrelated domains. Cross-domain comparisons and interpretations can be complex and cumbersome if the privacy evaluation method applied in a domain-specific work is equally niche. Therefore, a comparable evaluation method must work in various domains. Nevertheless, specialized metrics are by no means obsolete, as they can provide a unique perspective on particular problems. However, they should not be applied in a general scenario unless they are accompanied by a universal metric that allows for comparability.
4.2 Applicability
The success of a universal privacy evaluation metric is highly dependent on the applicability of the metric. A universal metric should be applicable to any synthetically generated tabular data, regardless of how it was generated and regardless of the data types. Furthermore, it should be realistic to compute the metric with available information and computational resources.
A1 Heterogeneity
The heterogeneous nature of tabular data can pose a challenge when evaluating privacy. Many evaluation methods rely on a measure of similarity between synthetic and true records. However, many popular similarity functions, such as Euclidean and Hamming distance, perform poorly with more than one specific data type. Commonly, authors declare that any similarity function can be applied and rely on the reader to implement the metric with necessary adjustments [20, 4]. Proposing metrics with vague definitions restricts the ability to compare results across various implementations accurately. Accordingly, a good privacy evaluation metric considers how data with heterogeneous data types should be handled.
A2 Diverse Generation Methods
The development and identification of generation methods and privacy implementations that balance utility and privacy is an important aspect of privacy-aware synthetic data research. Therefore, it is crucial that a universal privacy metric is applicable to diverse generation methods. Parametric generation methods imprint a certain characteristic on the synthetic data, which some evaluation methods can potentially favor. Furthermore, evaluation methods that are run during model training or fitting and are dependent on the iterative nature of such processes are likely to favor the methods for which they are designed. These evaluation methods negatively affect the applicability and, therefore, the comparability of privacy metrics.
A3 Practicality
Computational efficiency is highly relevant when considering how realistic the method is to use. Considering that many evaluation methods are based on pairwise distances [11, 14], the runtime should ideally be no more than $\mathcal{O}(n^{2})$ where $\mathcal{T},\mathcal{S}$ are the true and syntehtic data, respectively, and $n=\max(|\mathcal{T}|,|\mathcal{S}|)$.
A4 Implementation
A privacy evaluation method is only useful if it is realistic to implement and compute. Ideally, a well-documented implementation of the evaluation metric should be provided whenever a new method is proposed. At the very least, the method should be well documented to allow for easy implementation by the reader and to avoid any ambiguity. Consequently, a measure will gain a low implementation score if it contains many complex elements that require extensive domain knowledge to implement appropriately. Furthermore, the data required to compute the metric should be readily accessible, without external data being required. We distinguish between internal and external information: Internal information is contained in the datasets for which the metric is computed, and external information is gathered elsewhere. An example of this property is illustrated in Section 5.
4.3 Interpretability
An important reason for ensuring the privacy level in synthetic data is to enable data sharing of otherwise sensitive data. Regulatory agencies must recognize that adequate privacy levels are provided to accomplish that goal. Consequently, it is crucial that privacy levels can be accurately communicated to relevant stakeholders, be they data regulators or organizational managers. The Interpretability category aims to put emphasis on this aspect of creating universal privacy evaluation metrics.
I1 Explainability
What is measured and how it evaluates privacy should be easy to explain to any stakeholder. Using an explainable metric increases the transparency and, thereby, improves the confidence in the metric. Additionally, explainability increases the probability that the privacy provided by synthetic data is recognized and accepted by any relevant stakeholder. Accordingly, a good privacy measure is made up of elements that are simple to relay to laypeople.
I2 Understandability
An understandable metric is easy for laypeople to interpret without any additional explanation. That is, the output of the metric is understandable for laypeople. Naturally, a number without context is meaningless. Therefore, the criterion for an understandable metric is that it can be phrased in such a way that no ambiguity is present and no elaboration is necessary. An example could be “the risk of re-identification is 9%”, which would be more straightforward to interpret than something such as the evaluation of an adversarial prediction model “recall of 0.2 and precision of 0.3”.
I3 Visualization
Visualization is an important tool for improved understanding and reduced cognitive load. It allows people to see the data not only in isolation, but in comparison to one another. Therefore, a natural addition to Interpretability is visualization. Naturally, metrics that output single values can be visualized both in isolation and in comparisons, but so can other types of output whether it is multiple values, a distribution, or a range. Having visualization as a dimension allows metrics that produce multi-valued outputs to also receive a decent Interpretability score as long as they can be interpreted with the right means.
I4 Granularity
Granularity refers to a user’s ability to unambiguously differentiate between privacy levels. The most basic form of fine-grained differentiation is to ensure that the privacy measure follows a monotonic function. Any non-monotonic function can confuse a user, especially laypeople. Furthermore, a linear function is preferred as the privacy levels change at a constant rate as a function of the underlying computation. Consequently, any privacy metric that forces well-defined bounds by transformation, for example, by using a sigmoid or hyperbolic tangent function, will be slightly penalized. In addition, the level of granularity decreases if the output of the metric consists of more than a single value. Keeping track of more than one value means that the user must understand and interpret the individual outputs and how they change in connection. Lastly, the privacy level should be understandable in isolation. The results of some measures may require comparisons with the results of other datasets or generation methods to accurately differentiate between privacy levels. Such metrics should be penalized for this need.
4.4 Representativeness
The last category scores the ability of the privacy metric to accurately measure privacy levels. If the metric cannot capture potential privacy violations, the metric is not usable as a privacy metric.
R1 Anomalies
Anomalous records typically have a higher risk of being subject to privacy violations [20]. Inliers typically have the benefit of “blending in” with other common observations, ensuring that even highly similar synthetic records provide sufficient privacy levels. Outliers do not enjoy the same security. Synthetic records in the immediate proximity of a true record are likely to be uniquely linked to each other, constituting a privacy violation. Accordingly, anomalous observations should be given appropriate considerations when evaluating the overall level of privacy. Considering that many methods use some form of similarity measure, methods can be constructed to include weighted versions of the similarity function with the weights representing the degree to which individual records are outliers. The outcome of this is that records close to an outlier are made to be more similar to the outlier than would be the case if an unweighted distance function were used.
R2 Coverage
The privacy level reported by any metric should be accurate for all the records that the metric claims to include. Accordingly, if the privacy result is only accurate for some subsets of the data, the privacy measure does not represent the overall privacy level. One of the most obvious traps is to give too much weight to anomalous observations, thereby making the privacy level accurate only for outliers or even a subset of the outliers. This restriction is intended to create a balance between R1 and R2 such that it is not possible to obtain an inflated score by assigning excessive weights to outliers.
R3 Reproducibility
Metrics that provide different results when applied by different people or at different times carry some form of uncertainty about the metric’s accuracy. Therefore, an ideal privacy metric produces the same result when the input is the same. Most metrics that rely on some form of randomization will inevitably produce different results. Such metrics include, but are not limited to, prediction methods that require a random initialization state and methods that evaluate randomized batches of the data. Nevertheless, results can be somewhat consistent even with randomized components and, in such cases, the error should be clearly stated. Including an error does not constitute returning an extra value as part of the I4 Granularity dimension.
R4 Precision
A good evaluation metric neither over- nor underestimates the level of privacy provided by the synthetic data. This property is referred to as the precision of the metric. Overestimating privacy levels can be detrimental to the individuals contained in the data by providing a false sense of security. On the other hand, underestimation provides a conservative estimate of the privacy level, which does not harm individuals, but it does constitute an obstacle in optimizing the utility-privacy trade-off.
5 Application of CAIR
Figure 2: Radar plots of CAIR scores for the selected privacy evaluation metrics. Dimensions are indexed according to the system in Section 4 and separated into their respective categories denoted by the legend.
A benefit of having a scoring system to assess evaluation metrics is the ability to score existing metrics and rank them accordingly. As an illustrative example, a selection of tabular evaluation metrics is selected and scored with the results presented in Figure 2 (details on the scoring in Appendix A).
Distance to the closest record (DCR) is a simple metric where the output is typically the mean or median minimum distance from a synthetic record to a true one [9]. DCR can be used as a metric on its own or it can be extended to membership inference.
Threshold-based membership inference attacks (T. MIA) are attacks in which an adversary attempts to infer whether a known individual has been used to generate synthetic data [19]. Evaluation can be carried out in many ways, but a naïve approach is to use DCR for each record and to predict it as a member if $\text{DCR}\leq\tau$ where $\tau$ is a fixed threshold.
Attribute disclosure attacks (Attr. disclosure) are similar to MIAs, but instead of inferring membership, the adversary attempts to predict unknown attribute values [13]. A $k$-nearest neighbor algorithm can be used, and the majority vote predicts the univariate values for $k>1$. The attack is typically performed on various levels of prior adversarial knowledge and the output is all of the results or aggregate statistics.
$\epsilon$-identifiability ($\epsilon$-Ident.) takes a different approach and reports the proportion of synthetic records that are “too similar” to the true data points [20]. Let $d$ be the minimum weighted distance within the true data, let $\hat{d}$ be the weighted distance between the true and a specific synthetic record, and let $n$ be the number of synthetic records; then $\epsilon\text{-iden.}=\frac{1}{n}[\mathbb{I}(\hat{d}<d)]$, where $\mathbb{I}$ is the identity function.
Identity disclosure risk (IDR) incorporates functions like $\epsilon$-identifiability, but extends them with additional information, some of which are not readily available [4]. The function takes the inverse size of the equivalence class and multiplies by: 1) an indicator variable, such as $\epsilon$-identifiability, 2) an error correction term from the literature, and 3) a binary variable indicating whether the adversary learns something new. The equivalence class from both the sample and the true population is required.
The overall CAIR scores are given by the means with the standard error in Figure 2 where the best metric clearly is $\epsilon$-identifiability with a score of $3.56\pm 0.15$. The metric scores well in most dimensions and notably has some high scores in all four main categories. In contrast, the most closely related metric to $\epsilon$-identifiability, IDR, also performs well, but is especially penalized in the Applicability category for requiring information that is difficult to obtain such as equivalence classes of the true population. Consequently, IDR gets a considerably lower CAIR score, not because the metric is less accurate, but because it is difficult to employ and, thereby, not the best contender for a universal privacy metric.
Furthermore, DCR and attribute disclosure have relatively similar scores, but when investigating individual dimensions, it is evident that the two metrics have different downsides. Although DCR performs poorly on Comparability, the attribute disclosure metric has mainly poor scores for Interpretability. This exemplifies that CAIR not only allows stakeholders to rank metrics, but also allows researchers to accurately identify why they perform as they do and where improvements can be made. Interestingly, the same approach can be taken for well-performing metrics. $\epsilon$-identifiability can be improved in C3 data type agnostic by properly accommodating mixed data types and in R1 anomalies by improving the weighting of outliers.
In addition, threshold-based MIA and the attribute disclosure metric share many similarities, but their CAIR scores differ substantially. What is noticeable by investigating Figure 2 is that the Comparability dimensions are scored identically, while the largest difference is in Applicability and a few other dimensions. Based on the difference in Applicability observed in the figure, it can be estimated that attribute disclosure is somewhat more complex compared to threshold-based membership inference attacks.
Overall, these results suggest that the use of a rubric to break down the assessment of privacy evaluation metrics provides detailed insights into strengths and weaknesses of the metrics in comparison and isolation.
6 Discussion and Conclusion
With the increasing awareness of privacy-related issues in data management and sharing, it is more important than ever to focus on streamlining privacy evaluation, especially for maturing research fields such as synthetic data generation. Many domains, especially healthcare, rely heavily on the sharing of data to move the associated research fields forward. Unfortunately, the regulations put in place to protect people decelerate innovations. On the other hand, regulations force institutions to consider the natural right of privacy of people. Therefore, we must work with regulators to best ensure the continuation of potentially lifesaving innovation while still maintaining the privacy of the natural person.
CAIR is an important step in a direction in which regulators and researchers collaborate and enable data sharing. Regulators are starting to notice synthetic data as a possible substitute for true data [6]. Hence, a proper evaluation of the privacy-preserving capabilities of synthetic data is of great importance. A natural first step is to define what constitutes a good universal privacy-evaluation metric for synthetic tabular data. To the best of our knowledge, CAIR is the first attempt to define properties for such measures. Therefore, this work has the potential to guide the research community in a specific direction rather than using arbitrary metrics that researchers find appropriate for their work. Additionally, CAIR serves as guiding principles for regulators in identifying metrics and thresholds for when synthetic tabular data are “private enough” to be shared publicly. In summary, CAIR can help researchers, institutes, and regulators find common ground and agree on what universal privacy metrics are appropriate for synthetic tabular data.
One of the strengths of CAIR is the ability to rank metrics according to the 16 identified dimensions. However, CAIR does not guarantee that a high-scoring metric will be the most accurate metric for any given synthesis or synthetic data. This is not necessarily an issue with the definition of CAIR, but rather with the ideology of having a universal evaluation method. Therefore, we encourage to use universal metrics alongside metrics that are more specific for the given use case. In addition, the dimensions that have been identified as properties for good metrics are likely non-exhaustive. Consequently, future work should look at refining these properties while ensuring alignment with stakeholders. Furthermore, CAIR incentivizes researchers to consider what improvements can be made to existing privacy metrics and to design new metrics that conform to CAIR and the idea of universal evaluation metrics.